OpenStudy (anonymous):

I'm not sure how to do this, please don't just give the answer and show me how to do it step by step. Divide (9x^4y^3 + 3x^3y^2 - 6x^2y - 12x^2y^4) by -3x^2y

OpenStudy (phi):

term by term

OpenStudy (anonymous):

hey can you use equation editor and right this one again.

OpenStudy (anonymous):

Yes, of course, sorry.

OpenStudy (anonymous):

\[\frac{9x^4y^3}{-3x^2y}=?\]

OpenStudy (anonymous):

Divide \[(9x^4y^3 + 3x^3y^2 - 6x^2y - 12x^2y^4) \] by \[ -3x^2y\]

OpenStudy (anonymous):

Also @phi what do you mean "term by term"?

OpenStudy (anonymous):

like what i wrote, that is the first term

OpenStudy (anonymous):

Oh! Ah, sorry, I must have missed that when I was rewriting for newtonson. How would I be able to find the answer for the term? I haven't done this in a while, I don't remember much.

RhondaSommer (rhondasommer):

Welcome to Openstudy; Becayse Satellite73 helped you so much; a natrual thing to do is to hit the "Best Response" button. It is a sort of reward for those who help you :)

OpenStudy (anonymous):

I didn't really get any help though?? I was just told term by term and given an equation, I still have absolutely no clue what I'm supposed to be doing.

RhondaSommer (rhondasommer):

ah; then I shall help you

OpenStudy (anonymous):

Thank you. I would appreciate that.

RhondaSommer (rhondasommer):

:)

RhondaSommer (rhondasommer):

\[ (9x^4y^3 + 3x^3y^2 - 6x^2y - 12x^2y^4) /(-3x^2y)\]

RhondaSommer (rhondasommer):

you can simplify the x and y's by dividing\[x^2y\]. \[9x^4y^3/x^2y=9x^2y^2\]

RhondaSommer (rhondasommer):

you basically subtract the exponents

RhondaSommer (rhondasommer):

try that for \[3x^3y^2\]

OpenStudy (anonymous):

Would it just be 3x(y)?? No more exponents? Since when you subtracted 9x^4y^3, you ended up with 9x^2y^2.. I'm really sorry I'm not picking this up very well.

RhondaSommer (rhondasommer):

its okay :) \[3x^3y^2\] -----\[x^2y\] \[x^3/x^2~is~just~x^1\]

OpenStudy (anonymous):

Yes. I thought when you only had 1 exponent, they usually took the exponent off since it would just go back to being x anyways.

RhondaSommer (rhondasommer):

\[y^5/y^3=y^2\] and yes you are correct; i was just giving an example :)

RhondaSommer (rhondasommer):

examples*

OpenStudy (anonymous):

Oh, well I'm glad I at least remembered something for this!

OpenStudy (anonymous):

I think I know how to do this now, so thank you for helping!

RhondaSommer (rhondasommer):

yeah no problem :